# Autoregressive models for panel data causal inference with application to state-level opioid policies

### v1

@@ 2.2 Estimation
Here we describe approaches to estimating $\Delta_{t}$. We make one final assumption that facilitates the specification of estimators for the average treatment effect.
Assumption 6 (Time homogeneity).
For some $\Delta$, we have $\Delta_{t}=\Delta,\forall t.$
The time homogeneity assumption ensures that $\Delta_{t-1}$ (which we require for identification of $\Delta_{t}$) is identical to $\Delta_{t}$. We now review a few approaches to estimating $\Delta$. Under Assumptions 1-6, we can proceed by solving the following equation
$\psi_{1}=n^{-1}\sum_{i=1}^{n}Y_{it}\left[\frac{A_{it}}{e(Y_{i,t-1}-A_{i,t-1}\Delta)}-\frac{1-A_{it}}{1-e(Y_{i,t-1}-A_{i,t-1}\Delta)}\right]-\Delta=0.$
We cannot directly solve this without specification of $e(\cdot)$, so one could specify a parametric model $g\left\{e(Y_{i,t-1};\Delta)\right\}=\beta_{0}+\beta_{1}\left(Y_{i,t-1}-A_{i,t-1}\Delta\right)$ for a known link function $g(\cdot)$, and these two sets of parameters can be solved jointly using M-estimation theory.
@@
Alternatively, we can specify a model for the conditional expectation function in equation (1), such as a debiased linear autoregressive model (Benjamin, Rigby and Stasinopoulos, 2003; Zeger and Qaqish, 1988) (L-DAM):
$Y_{it}$ $=\alpha+\delta(Y_{i,t-1}-\Delta A_{i,t-1})+\Delta A_{it}+\epsilon_{it}$ (2)
@@
where we assume that the errors $\epsilon_{it}$ are conditionally mean zero with zero correlation over time. The L-DAM is essentially a simple linear parameterization of the causal assumptions presented above. We may once again use maximum likelihood to obtain estimates and inference for this model using M-estimation. Alternatively, we may specify priors for the parameters in (2) and use Bayesian inference, the approach we adopt in the rest of the paper.
These approaches could easily be adapted if Assumption 6 is implausible, or if time-specific effects are of particular interest. However, Assumption 6 connects our estimation approach to the traditional L-DAM in (2). Also, in many studies of state policies, at any given time, because data are limited, making assumptions like Assumption 6 that reduce the number of parameters to estimate may strike a favorable bias-variance trade-off.
2.3 Impacts of treatment effect heterogeneity
Assumption 4, which enforces constant treatment effects across units, is an especially strong assumption required for our estimation strategy. Under our modeling assumptions, it requires that $Y_{i,t-1}-\Delta A_{i,t-1}$ is equal to $Y_{i,t-1}(0)$. In the presence of treatment effect heterogeneity, this equivalence will not hold, leading standard estimation strategies to have both biased effect and variance estimates. Moreover, as heterogeneity increases, the degree of this bias increases. Note that there are many ways that ignoring treatment effect heterogeneity can negatively impact results that are not unique to our setting. However, because this particular form of misspecification – where we use estimated proxies to control for confounders – is unique to our identification strategy, we focus our discussion in this section on the impact of this misspecification, particularly with respect to the bias of the effect estimates.
2.3.1 Treatment effect heterogeneity as measurement error
@@
Suppose Assumption 4 does not hold. Define $\Delta_{it}=Y_{it}(1)-Y_{it}(0)$ as the unit-level treatment effect for unit $i$ at time $t$ such that $\Delta=\mathbb{E}(\Delta_{it})$. We can re-write (2) to obtain
$Y_{it}$ $=\alpha+\delta(Y_{i,t-1}(0)+A_{i,t-1}\Delta_{i,t-1}-A_{i,t-1}\Delta)+\Delta A_{it}+\epsilon_{it}$
$=\alpha+\delta(Y_{i,t-1}(0)+U_{i,t-1})+\Delta A_{it}+\epsilon_{it}.$ (3)
@@
In other words, adjusting for $Y_{i,t-1}-A_{i,t-1}\Delta$, our proxy of $Y_{i,t-1}(0)$, can be equivalently thought of as adjusting for a noisy version of $Y_{i,t-1}(0)$ subject to measurement error. An important parameter therefore becomes $\text{Var}(U_{i,t-1})$, which controls the magnitude of the heterogeneity and measurement error. Clearly, if $\text{Var}(U_{i,t-1})=0$ and we have no heterogeneity, then Assumption 4 holds, and Theorem 1 states that we can identify $\Delta$.
@@
Throughout the rest of this section, we focus on a simplified $\widehat{\Delta}$ that comes from estimating the model in (3), ignoring the fact that $U_{i,t-1}$ is a function of $\Delta$, and focusing on the impact of this sort of measurement error in this setting. [^1] We find identical results when fitting the model in (2), so we proceed with the simpler setting that has easier to understand mathematical derivations. Measurement error in a confounder has been studied previously (Carroll, 1998), and it was shown that if $U_{i,t-1}$ is mutually independent of $Y_{i,t-1}(0),A_{it}$, and $Y_{it}$, then
@@
$\mathbb{E}(\widehat{\Delta})=\Delta+\delta\gamma_{A}\bigg{(}\frac{\text{Var}(U_{i,t-1})}{\text{Var}(U_{i,t-1})+\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}}\bigg{)},$ (4)
@@
where $\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}$ is the residual variance from regressing $Y_{i,t-1}(0)$ against $A_{it}$, and $\gamma_{A}$ is the coefficient for $A_{it}$ in that regression. If we did not adjust for $Y_{i,t-1}(0)$ at all, we would have $\mathbb{E}(\widehat{\Delta})=\Delta+\delta\gamma_{A}$. This shows that adjusting for an error prone version of $Y_{i,t-1}(0)$ (instead of $Y_{i,t-1}(0)$ itself) biases estimates towards what we would obtain without adjusting for the confounder at all, and the degree of this bias depends on the relative size of $\text{Var}(U_{i,t-1})$ and $\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}$. [^2] In state-level studies of opioid policies, we expect the treatment effect $\Delta$ to be relatively small in magnitude relative to the overall degree of variation in the outcomes, leading this bias to be small. Proposition 1 formalizes reasoning that may be used to justify the proposed estimator.
@@
Under Assumptions 1, 3, 5, 6 and assuming model (3) holds, if $U_{i,t-1}$ is mutually independent of $Y_{i,t-1}(0),A_{it}$, and $Y_{it}$, and there exists a $k>0$ such that $\frac{\text{Var}(U_{i,t-1})}{\sigma_{Y_{i,t-1}(0)\mid A_{it}}}<k$, then we have
$\left|\mathbb{E}(\widehat{\Delta})-\Delta\right|\leq|\delta\gamma_{A}|\frac{k}{k+1}$
Proposition 1 shows that only large amounts of heterogeneity would lead to substantial bias in treatment effect estimates when the measurement error follows a classical structure and is independent of the treatment, outcome, and potential outcomes. However, this may not hold in many settings. We therefore next discuss the impacts of differential measurement error.
@@ 2.3.2 Differential measurement error and sources of heterogeneity
If the measurement error terms $U_{i,t-1}$ are correlated with any of $A_{it}$, $Y_{it}$, or $Y_{it}(0)$ then the bias expression in (4) no longer holds. The exact bias expression in this setting can be found in Appendix B, where we discuss the impacts of differential measurement error from treatment effect heterogeneity in greater detail. While we leave most details to the Appendix, we focus here on a particular form of differential measurement error that is most likely to occur and is most problematic.
The most likely reason in our context for differential measurement error arises due to correlation between $U_{i,t-1}$ and $Y_{it}$. When $U_{i,t-1}$ is high, unit $i$ was treated and had a large treatment effect at time $t-1$. When unit-level treatment effects are smooth over time, a large value of $U_{i,t-1}$ likely correlates with a high value of $Y_{it}$ due to the large unit-level treatment effect. Applying the formula from Appendix B to the situation where $U_{i,t-1}$ is correlated with $Y_{it}$, but not $A_{it}$, the estimate of the treatment effect has expectation
$\mathbb{E}(\widehat{\Delta})$ $=\frac{C_{1}+C_{2}+C_{3}}{\text{Var}(A_{it})-\frac{\text{Cov}^{2}(A_{it},Y_{i,t-1}(0))}{\text{Var}(Y_{i,t-1}(0)+U_{i,t-1})}}$
$=\Delta+\delta\gamma_{A}\bigg{(}\frac{\text{Var}(U_{i,t-1})}{\text{Var}(U_{i,t-1})+\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}}\bigg{)}+\frac{C_{3}}{\text{Var}(A_{it})-\frac{\text{Cov}^{2}(A_{it},Y_{i,t-1}(0))}{\text{Var}(Y_{i,t-1}(0)+U_{i,t-1})}},$
@@
Notice that when the measurement error is classical and $\text{Cov}(Y_{it},U_{i,t-1})=0$, then $C_{3}=0$ and we revert back to the value in (4). Now we can examine whether $C_{3}$ is small (or large) in magnitude compared with $C_{1}$ and $C_{2}$ to examine whether the fact that the measurement error is differential plays a significant role in the magnitude of bias. We can see that the ratio of $C_{3}$ and $C_{2}$ can be written as,
@@
2.3.3 Addressing treatment effect heterogeneity
We propose two strategies to proceed when estimating treatment effects using the autoregressive model (equation 2) in the presence of effect heterogeneity: in short, we can either ignore or model the heterogeneity. In most instances, it will be reasonable to assume that the hetereogeneity, and therefore the induced bias, is small. We explore this further in Section 5, where we find that heterogeneity of the treatment effect only impacts inference on the average treatment effect when the degree of heterogeneity is very large.
For cases where we cannot assume that effect heterogeneity is sufficiently small, we can modify the autoregressive framework to model the effect heterogeneity. For example, if $\Delta_{it}=h(X_{i},t)$, where $X_{i}$ are observed covariates, then we can easily specify a parametric form for $h(\cdot)$. We can then adjust for $Y_{i,t-1}-A_{i,t-1}h(X_{i},t-1)$, which should be a good proxy for the true $Y_{i,t-1}(0)$ for a reasonably well-specified $h(\cdot)$ function, alleviating effect heterogeneity induced biases. However, in some applications, we may have little data to estimate treatment effects, and this additional modeling may create too many parameters and lead to overly variable effect estimates. By contrast, in settings where we have relatively large sample sizes, this may be a perfectly reasonable strategy. On the other hand, if the heterogeneity is not a function of $X_{i}$, then we cannot use this approach and may instead rely on the results of the previous sections that argue this bias is likely small in magnitude. For example, if the heterogeneity takes the form $\Delta_{it}=h(V_{i})$ for missing variables $V_{i}$, then the resulting measurement error will likely be correlated with $Y_{it}$, as previously discussed.
One aspect unique to panel data that is crucial for this discussion is that we can partially assess the degree of heterogeneity empirically, and therefore whether the assumption that it does not significantly bias our results is reasonable. For treated unit-time pairs, we observe $Y_{it}=Y_{it}(0)+\Delta_{it}$. Assuming that the individual treatment effects $\Delta_{it}$ are not strongly and negatively correlated with $Y_{it}(0)$, then we would expect $\text{Var}(Y_{it}(0))\leq\text{Var}(Y_{it})$ for treated units at time $t$. We can estimate the variance of $Y_{it}$ during treated time periods and compare that with the variance of $Y_{it}$ during control periods, possibly adjusting for time-varying trends in variability. Similar ideas have been applied in randomized trials to test for heterogeneity (Mills et al., 2021). The idea is that a larger estimated outcome variance in the treated time periods would suggest heterogeneous treatment effects and would therefore potentially be a larger concern. This is not a formal test of whether heterogeneity is present, because these variances can be the same even in the presence of heterogeneity when $Y_{it}(0)$ and $\Delta_{it}$ are negatively correlated; nevertheless, this provides some additional information about the plausibility of the constant treatment effects assumption.
3 Relaxing assumptions and extensions
Up to this point we have focused on a simple set of identification assumptions to provide a straightforward and intuitive explanation of our identification and estimation strategies. While useful for exposition, these assumptions were quite strong, and can be weakened substantially. In this section, we relax these assumptions and show how these relaxations affect our proposed estimation strategy.

### v2

@@ 3.2 Estimation
Under Assumptions 1-6, we specify a debiased linear autoregression model (L-DAM) for the conditional expectation of the outcome (Benjamin, Rigby and Stasinopoulos, 2003; Zeger and Qaqish, 1988):
$Y_{it}$ $=\alpha+\beta(Y_{i,t-1}-\theta A_{i,t-1})+\theta A_{it}+\epsilon_{it}$ (3)
@@
where $\mathbb{E}[\epsilon_{it}\mid A_{it},Y_{i,t-1}(0)]=0$, and we assume that the errors have zero correlation across units or over time (i.e. $Cov(\epsilon_{it},\epsilon_{js})=0$ for $j\neq i,t\neq s$). The L-DAM reflects a simple linear parameterization of the assumptions above. We may use M-estimation to estimate these parameters and to conduct inference. Alternatively, we may specify priors for the parameters in (3) and use Bayesian inference, the approach we adopt in the rest of the paper. We also suggest an IPW estimator and a doubly-robust approach that consistently estimates $\theta$ if either the IPW or outcome model are misspecified in Appendix H.
4 Relaxing assumptions and extensions
We have focused on a simple set of identification assumptions to provide a straightforward and intuitive motivation for a simplified version of our proposed estimation strategy. While useful for exposition, these assumptions were quite strong and may be weakened substantially. We now consider both relaxations of these assumptions, and how violations of the assumptions may bias our estimates. Specifically, for this latter piece we focus our discussion on Assumption 4, which enforces constant treatment effects. While ignoring treatment effect heterogeneity can negatively impact results in many ways that are not unique to our setting, the particular form of misspecification we study is unique to our identification strategy, and we therefore focus much of the discussion on this problem, including a mathematical argument that ignoring this heterogeneity will not generally induce large biases. Finally, we consider extensions that relax other assumptions, such as allowing for additional time dependence, more complex estimands allowing for different treatment sequences, time-varying covariates, multiple policies, and non-linear outcome models.
4.1 Effect heterogeneity
Assumption 4 is an especially strong assumption required for identification and estimation. Under our modeling assumptions, it requires that $Y_{i,t-1}-\theta A_{i,t-1}$ is equal to $Y_{i,t-1}(0)$. In the presence of treatment effect heterogeneity, this equivalence will not hold, leading standard estimation strategies to have both biased effect and variance estimates. Moreover, as heterogeneity increases, the degree of this bias increases. We first discuss a simple solution to this problem: modeling the effect heterogeneity as a function of observed covariates. While a simple solution, we argue that in many applications this won’t be necessary, and later use mathematical derivations to argue that the bias due to effect heterogeneity is likely small in many settings.
4.1.1 Modeling the effect heterogeneity
A simple approach to allow effect heterogeneity is to model it within the autoregressive framework. For example, if $\mathbb{E}[Y_{it}(1)-Y_{it}(0)\mid V_{it}]=h(V_{it},t)$, where $V_{it}$ includes some subset of the observed covariates $X_{it}$, we can specify a parametric form for $h(\cdot)$ and adjust for $Y_{i,t-1}-A_{i,t-1}h(V_{i,t-1},t-1)$. [^2] It is also potentially feasible to allow heterogeneity with the prior period control counterfactuals. However, this would lead to a recursive formula requiring a time period where no one was treated in order to estimate the equation. A well-specified $h(\cdot)$ function will eliminate effect heterogeneity induced biases and allow us to study whether certain characteristics modify the treatment effect. However, one practical issue, particularly when analyzing state-level data, is that we already have little data to estimate treatment effects. This additional modeling may lead to overly variable effect estimates, as well as reduce power to estimate average effects.
We therefore propose an approach to estimating heterogeneous treatment effects through a parametric model for $h(V_{it},t)$ that can simultaneously identify treatment effect heterogeneity when it is present, while also reducing the impact of estimation uncertainty when the effects are at least approximately homogeneous or we have insufficient data to fit such models. Specifically, we assume the additive model
$h(V_{it},t)=\nu_{0}+b(t)+\sum_{j}V_{it,j}\nu_{j},$
where $b(\cdot)$ can be fit using any one-dimensional smoothing approach, though we focus our attention on estimation of the $\nu_{j}$ coefficients and heterogeneity by covariates rather than heterogeneity over time. Note that while we have included covariates $V_{it}$ into this model linearly, nonlinear functions of these variables could also be incorporated in a similar manner. Placing non-informative prior distributions on the $\nu_{j}$ coefficients might be the preferred strategy if we had large amounts of data; however, this will lead to overly variable estimates in small data settings or when the number of covariates included in $V_{it}$ is large. To reflect the fact that treatment effects both 1) may indeed be homogeneous, or at least approximately so, and 2) we may not have enough data to estimate these effects, we incorporate recent prior distributions in the high-dimensional Bayesian statistics literature. We can use standard, non-informative Gaussian prior distributions for $\nu_{0}$ and the coefficients of $b(\cdot)$ as we do not want to shrink the overall treatment effect towards zero. For the $\nu_{j}$ parameters that dictate heterogeneity by observed covariates or prior counterfactuals, we can use the horseshoe prior distribution (Carvalho, Polson and Scott, 2009) to shrink this heterogeneous treatment effect function towards a homogeneous effect. This prior distribution takes the form
$\nu_{j}$ $\sim\mathcal{N}(0,\upsilon\omega_{j})$
$\upsilon$ $\sim C^{+}(0,1)$
$\omega_{j}$ $\sim C^{+}(0,1)\quad\text{for all }j,$
where $C^{+}(0,1)$ represents the half-Cauchy prior distribution. This prior distribution has been shown to aggressively shrink coefficients corresponding to non-important predictors towards zero, while reducing bias in the estimation of large coefficients. This allows us to accommodate heterogeneity without sacrificing too much efficiency, which is critical in state-level policy applications. Additionally, we can use the posterior distribution of $\bm{\nu}$ to test the null hypothesis that there is no heterogeneity present. This null hypothesis corresponds to $\nu_{j}=0$ for all $j$, and we may construct simultaneous posterior credible intervals for all $\nu_{j}$ to address multiplicity concerns and examine whether zero is contained in the posterior distribution. Moreover, in panel data settings, there are potentially other tests one could construct that are not unique to the statistical model presented here. Roughly speaking, if treatment effect heterogeneity is present, we would expect more variability in treated units than control units. Similar ideas were explored in the context of randomized trials in Mills et al. (2021). Extending these to the observational, panel data setting here is beyond the scope of this paper, but is a potential topic for future research.
Of course, this strategy only works if the heterogeneity is a function of $V_{it}$. When it is not, this approach may still result in biased effect estimates for average causal effects. Fortunately, as we show formally below, in many instances it may be reasonable to assume that the heterogeneity, and therefore the induced bias, is small – even when we do not model the heterogeneity. We also explore this in our simulation study in Section 6, where we show that heterogeneity of the treatment effect only impacts inference on the average treatment effect when the degree of heterogeneity is very large. Ignoring this heterogeneity may therefore often be a reasonable option in practice, particularly in situations where the average treatment effect is of primary substantive interest.
4.1.2 Effect heterogeneity as measurement error
@@
We now derive an expression for the bias due to effect heterogeneity when constant treatment effects are assumed. We do this by reframing this as a measurement error problem. Suppose Assumption 4 does not hold. Define $\theta_{it}=Y_{it}(1)-Y_{it}(0)$ as the unit-level treatment effect for unit $i$ at time $t$ such that $\theta=\mathbb{E}(\theta_{it})$. We can re-write (3) to obtain
$Y_{it}$ $=\alpha+\beta(Y_{i,t-1}(0)+A_{i,t-1}\theta_{i,t-1}-A_{i,t-1}\theta)+\theta A_{it}+\epsilon_{it}$
$=\alpha+\beta(Y_{i,t-1}(0)+U_{i,t-1})+\theta A_{it}+\epsilon_{it}.$ (4)
@@
In other words, adjusting for $Y_{i,t-1}-A_{i,t-1}\theta$, our proxy of $Y_{i,t-1}(0)$, can be equivalently thought of as adjusting for a noisy version of $Y_{i,t-1}(0)$ subject to measurement error. An important parameter therefore becomes $\text{Var}(U_{i,t-1})$, which controls the magnitude of the heterogeneity and measurement error. Clearly, if $\text{Var}(U_{i,t-1})=0$ and we have no heterogeneity, then Assumption 4 holds, and $\theta$ is identified.
@@
Throughout the rest of this section, we focus on a simplified $\widehat{\theta}$ that comes from estimating the model in (4), ignoring the fact that $U_{i,t-1}$ is a function of $\theta$, and focusing on the impact of this sort of measurement error in this setting. [^3] We find identical results when fitting the model in (3), so we proceed with the simpler setting that has easier to understand mathematical derivations. Measurement error in a confounder has been studied previously (Carroll, 1998), and it was shown that if $U_{i,t-1}$ is mutually independent of $Y_{i,t-1}(0),A_{it}$, and $Y_{it}$, then
@@
$\mathbb{E}(\widehat{\theta})=\theta+\beta\gamma_{A}\bigg(\frac{\text{Var}(U_{i,t-1})}{\text{Var}(U_{i,t-1})+\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}}\bigg),$ (5)
@@
where $\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}$ is the residual variance from regressing $Y_{i,t-1}(0)$ against $A_{it}$, and $\gamma_{A}$ is the coefficient for $A_{it}$ in this regression. If we did not adjust for $Y_{i,t-1}(0)$ at all, we would have $\mathbb{E}(\widehat{\theta})-\theta=\beta\gamma_{A}$. In other words, adjusting for an error-prone version of $Y_{i,t-1}(0)$ instead of $Y_{i,t-1}(0)$ itself biases estimates towards what we would obtain without adjusting for the confounder at all. Moreover, the degree of this bias depends on the relative size of $\text{Var}(U_{i,t-1})$ and $\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}$. [^4] In state-level studies of opioid policies, we expect the treatment effect $\theta$ to be relatively small in magnitude relative to the overall degree of variation in the outcomes, leading this bias to be small. Proposition 3 formalizes reasoning that may be used to justify the proposed estimator.
@@
Under Assumptions 1, 2, 3, 5 and assuming model (4) holds, if $U_{i,t-1}$ is mutually independent of $Y_{i,t-1}(0),A_{it}$, and $Y_{it}$, and there exists a $k>0$ such that $\frac{\text{Var}(U_{i,t-1})}{\sigma_{Y_{i,t-1}(0)\mid A_{it}}}<k$, then we have
$\left|\mathbb{E}(\widehat{\theta})-\theta\right|\leq|\beta\gamma_{A}|\frac{k}{k+1}$
Proposition 3 shows that only large amounts of heterogeneity would lead to substantial bias in the estimation of the treatment effect when the measurement error follows a classical structure and is independent of the treatment, outcome and potential outcomes. However, the classical structure may not hold in many settings. Next, we discuss the impacts of differential measurement error.
@@ 4.1.3 Differential measurement error and sources of heterogeneity
If the measurement error terms $U_{i,t-1}$ are correlated with any of $A_{it}$, $Y_{it}$, or $Y_{it}(0)$ then the bias expression in (5) no longer holds. The exact bias expression in this setting can be found in Appendix B, where we discuss the impacts of differential measurement error from treatment effect heterogeneity in more detail. While we leave most details to the Appendix, we focus here on a particular form of differential measurement error that is most likely to occur and is most problematic.
The most likely reason in our context for differential measurement error arises from correlation between $U_{i,t-1}$ and $Y_{it}$. When $U_{i,t-1}$ is high, unit $i$ was treated and had a large treatment effect at time t1t-1. When unit-level treatment effects are smooth over time, a large value of $U_{i,t-1}$ likely correlates with a high value of $Y_{it}$ due to the large unit-level treatment effect. Applying the formula from Appendix B to the situation where $U_{i,t-1}$ is correlated with $Y_{it}$, but not $A_{it}$, the estimate of the treatment effect has expectation
[…]
$=\theta+\beta\gamma_{A}\bigg(\frac{\text{Var}(U_{i,t-1})}{\text{Var}(U_{i,t-1})+\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}}\bigg)+\frac{C_{3}}{\text{Var}(A_{it})-\frac{\text{Cov}^{2}(A_{it},Y_{i,t-1}(0))}{\text{Var}(Y_{i,t-1}(0)+U_{i,t-1})}},$
@@
Notice that when the measurement error is classical and $\text{Cov}(Y_{it},U_{i,t-1})=0$, then $C_{3}=0$ and we revert back to the value in (5).
We can examine whether $C_{3}$ is small (or large) in magnitude compared with $C_{1}$ and $C_{2}$ to assess how differential measurement error affects the magnitude of bias. We can see that the ratio of $C_{3}$ and $C_{2}$ can be written as,
@@
In summary, the mathematical expressions for the bias induced by effect heterogeneity when constant treatment effects are assumed suggest that for many practical applications, we may reasonably assume this bias to be small.